x+5z=0
5x+y=0
y-3z=0
Help! Solve

Answer by TimothyLamb(4379)

---
x+5z=0
5x+y=0
y-3z=0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
solution:
x = 0
y = 0
z = 0
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php